alibaby4562:

if x^2 = 100 what is x?

Buenos88:

\[\sqrt{100}=x\]

MrMudd183:

that's the most accurate way to get the answer. since 100 is a perfect square, then the answer would be 10. and 10 can be squared to 100.

alibaby4562:

Use the square root property to solve the equation. The equation has real number solutions.

ramen:

what times itself multiplies to 100

Buenos88:

\[\pm10\] meaning -10 and 10 could be x.

TheSmartOne:

@mrmudd183 You have to remember that the square root will result in \(\pm\) values

Zarkon:

\(x^2=100\) \(x^2-100=0\) \((x-10)(x+10)=0\) so \(x=10\) or \(x=-10\)

MrMudd183:

ok, thank you TSO. i kind of forgot about that. it's kind of been a while. but i'm trying my best to help, but math isn't my strongest subject

ramen:

3 ways: sqrt100, what times it self = 100 (mental maths), or zarkon's way